OpenStudy (anonymous):

Please help me set up an equation! Your piggy bank has a total of 43 coins in it; some are dimes and some are nickels. If you have a total of $3.60, how many dimes do you have? Once I have an equation I'm pretty sure I will be able to solve it, I just need help making the equation

OpenStudy (johnweldon1993):

Well we know there are dimes and nickles that make up those 43 coins...so \[\large D + N = 43\] We know that a dime is .10 and a nickle is .05....and the total is 3.60...so \[\large .10D + .05N = 3.60\]

OpenStudy (anonymous):

okay! Thank you!

OpenStudy (johnweldon1993):

No problem!

OpenStudy (anonymous):

Sometimes calculating in pennies may be easier.\[\{d+n=43,10d+5n=360\}\]\[\{d=29,n=14\} \]

OpenStudy (johnweldon1993):

Yeah thats true too lol...if you don't like working with decimals....multiply the second equation by 100 to make whole numbers lol